# Bayesian Sheaf Neural Networks

### v1

@@ Appendix B Section 4 Proofs
Equip $SO(n)$ with the metric induced by the Frobenius norm $\|A\|=\operatorname{tr}(A^{\mathsf{T}}A)$. Since left (or right) multiplication on $SO(n)$ is an isometry, it follows that the Hausdorff measure is a left (and right) translation invariant measure on $SO(n)$. Moreover, if $d=n(n-1)/2$ denotes the dimension of $SO(n)$, then $\mathcal{H}^{d}$ is a locally finite measure on $SO(n)$, and thus is a Haar measure by Haar’s theorem.
Hence, Proposition B.1 yields a change of variables formula for they Cayley transform. For example, if $X$ is a uniformly distributed random variable on $SO(n)$, then $C^{-1}(X)$ has density $p(x)=J\widetilde{C}(x)$ with respect to the Lebesgue measure on $\mathbb{R}^{d}$.
Proof of Theorem 4.4.
[…]
@@
Proof of Proposition 4.6.
Representing points of $S^{3}$ as unit vectors $x=(a,b,c,d)\in\mathbb{R}^{4}$, the $2$--fold cover $\Phi:S^{3}\to SO(3)$ has an explicit description as
@@
Appendix C KL Divergence Calculations
Recall that we assume that the variational distribution $q_{\phi}(\mathcal{F}|X,y)$ factors as $q_{\phi}(\mathcal{F}|X,y)=\prod_{u\trianglelefteq e}q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)$ over all incident node-edge pairs $u\trianglelefteq e$, and that we use a prior $p_{\theta}(\mathcal{F})$ which also factors in the same manner $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$. As a consequence of the chain rule for KL divergence, see for example [17, Theorem 2.5.3], we have
@@ Appendix C KL Divergence Calculations
In the cases where we learn diagonal or general linear sheaf restriction maps, $\textnormal{KL}(q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)\;\|\;p_{\theta}(\mathcal{F}_{u\trianglelefteq e}))$ is a KL divergence of a normal distribution with diagonal covariance from a standard normal distribution. The KL divergence of a normal distribution $\mathcal{N}(\mu,\operatorname{diag}(\sigma^{2}))$ with mean $\mu\in\mathbb{R}^{d}$ and $\sigma^{2}\in\mathbb{R}^{d}$ from the standard normal distribution is
@@
In the case where the restriction maps are special orthogonal, and the stalk dimension $d$ is $2$ or $3$, we compute $\textnormal{KL}(q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)\;\|\;p_{\theta}(\mathcal{F}_{u\trianglelefteq e}))$ using Proposition 4.7, which we now prove.
Proof of Proposition 4.7.
@@
To simplify notation, set $d=n(n-1)/2$, let $\gamma=\frac{1+\kappa}{1-\kappa}$, and let $V_{n}$ denote the volume of $SO(n)$. Recall Eq. (17), and note that the density of $\mathcal{C}_{n}(I,\kappa)$ with respect to $\mathcal{H}^{d}$ can be expressed as
$V_{n}^{-1}\gamma^{d}\Big{(}\frac{\det(I+C^{-1}(P)}{\det(I+\gamma C^{-1}(P))}\Big{)}^{n-1}$
where $V_{n}$ denotes the volume of $SO(n)$. Then the change of variables formula yields
@@
where recall from Section 4 that $X_{\phi}$ denotes the skew symmetric matrix whose entries below the diagonal are the entries of the vector $\phi$.
@@
$\textnormal{KL}(\operatorname{Cauchy}_{1}(0,\tfrac{1}{\gamma^{2}}I)\;\|\;\operatorname{Cauchy}_{1}(0,I))=\log\bigg{(}\frac{(\gamma+1)^{2}}{4\gamma}\bigg{)}$
@@
Appendix D Experimental Details
All experiments were run on a NVIDIA Tesla K80 GPU with two Intel Xeon E5-2637 CPUs with 64GB of memory. Hyperparameter tuning was done using Weights & Biases [4] to perform a grid search over the hyperparameter configurations described in Table 3. The neural sheaf diffusion (NSD) models from [6] were implemented using the code available at https://github.com/twitter-research/neural-sheaf-diffusion. Our code is available at https://github.com/patrick-gillespie/bsnn.
Table 3: Hyperparameter configurations for WebKB experiments.
@@ Appendix D Experimental Details
D.1 WebKB Dataset Descriptions
The Texas, Cornell, and Wisconsin datasets are web page datasets for the computer science departments of the respective universities. Each dataset is a graph in which nodes represent web pages and edges represent hyperlinks between them. In all datasets the node features are $1703$--dimensional bag-of-word representations of the web pages, and each node belongs to one of the five classes: student, project, course, staff, and faculty. The Texas dataset contains 183 nodes and 325 edges. The Cornell dataset contains 183 nodes and 298 edges. The Wisconsin dataset contains 251 nodes and 515 edges.

### v2

@@ Acknowledgments
Equip $SO(n)$ with the metric induced by the Frobenius norm $\|A\|=\sqrt{\operatorname{tr}(A^{\mathsf{T}}A)}$. Since left (or right) multiplication on $SO(n)$ is an isometry, it follows that the Hausdorff measure is a left (and right) translation-invariant measure on $SO(n)$. Moreover, if d=n(n1)/2d=n(n-1)/2 denotes the dimension of $SO(n)$, then $\mathcal{H}^{d}$ is a locally finite measure on $SO(n)$, and thus is a Haar measure by Haar’s theorem.
Hence, Proposition 4 yields a change of variables formula for the Cayley transform. For example, if $X$ is a uniformly distributed random variable on $SO(n)$, then $C^{-1}(X)$ has density $p(x)=J\widetilde{C}(x)$ with respect to the Lebesgue measure on $\mathbb{R}^{d}$.
Proof of Theorem 1.
Let $\mu$ denote the unique Haar measure on $SO(n)$ such that $\int_{SO(n)}d\mu=1$. Let $X$ be a uniformly distributed random variable on $SO(n)$. Let $Y=C(\frac{1-\kappa}{1+\kappa}C^{-1}(X))$, and let $f_{n}(P;I,\kappa)$ denote the density of $Y$ with respect to $\mu$. For convenience, set $\gamma=\frac{1+\kappa}{1-\kappa}$. Then by Proposition 4 and the chain rule,
@@
Proof of Proposition 2.
Representing the points of $S^{3}$ as unit vectors $x=(a,b,c,d)\in\mathbb{R}^{4}$, the 22-fold cover $\Phi:S^{3}\to SO(3)$ has an explicit description as
@@
Appendix C: KL Divergence Calculations
Recall that we assume that the variational distribution $q_{\phi}(\mathcal{F}|X,y)$ factors as $q_{\phi}(\mathcal{F}|X,y)=\prod_{u\trianglelefteq e}q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)$ over all incident node-edge pairs $u\trianglelefteq e$, and that we use a prior $p_{\theta}(\mathcal{F})$ which also factors in the same manner $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$. As a consequence of the chain rule for KL divergence, see for example Theorem 2.5.3 in [22], we have
@@
In the cases where the sheaf restriction maps are diagonal or general linear, this divergence reduces to the KL between a Gaussian with diagonal covariance and a standard normal distribution. The KL divergence of a normal distribution $\mathcal{N}(\mu,\operatorname{diag}(\sigma^{2}))$ with mean $\mu\in\mathbb{R}^{d}$ and $\sigma^{2}\in\mathbb{R}^{d}$ from the standard normal distribution is
@@
In the case where the restriction maps are special orthogonal, and the stalk dimension $d$ is $2$ or $3$, we compute $\textnormal{KL}(q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)\;\|\;p_{\theta}(\mathcal{F}_{u\trianglelefteq e}))$ using Proposition 3, which we now prove.
Proof of Proposition 3.
@@
To simplify notation, set d=n(n1)/2d=n(n-1)/2, let $\gamma=\frac{1+\kappa}{1-\kappa}$, and let $V_{n}$ denote the volume of $SO(n)$. Recall Eq.(24), and note that the density of $\mathcal{C}_{n}(I,\kappa)$ with respect to $\mathcal{H}^{d}$ can be expressed as
$V_{n}^{-1}\gamma^{d}\Big{(}\frac{\det(I+C^{-1}(P)}{\det(I+\gamma C^{-1}(P))}\Big{)}^{n-1},$
where $V_{n}$ denotes the volume of $SO(n)$. Then the change of variables formula Eq.(23) yields
@@
where recall from Section 3.3 that $X_{\phi}$ denotes the skew symmetric matrix whose entries below the diagonal are the entries of the vector $\phi$.
@@
$\textnormal{KL}(\operatorname{Cauchy}_{1}(0,\tfrac{1}{\gamma^{2}}I)\;\|\;\operatorname{Cauchy}_{1}(0,I))=\log\bigg{(}\frac{(\gamma+1)^{2}}{4\gamma}\bigg{)},$
@@
Appendix D: Experiment Details
We selected a diverse set of real-world datasets exhibiting varying levels of homophily. Table 6 shows the number of nodes, edges, and classes in each dataset.
The Texas, Cornell, and Wisconsin datasets are part of the WebKB collection, each representing the computer science department of a respective university. Each dataset is a graph in which nodes represent web pages and edges represent hyperlinks between them. In all datasets the node features are 17031703-dimensional bag-of-word representations of the web pages.
The Film dataset is a graph of actors, where nodes represent actors and edges represent co-appearances in films. Node features are derived from metadata such as keywords and genres, and each node belongs to one of several genre-based classes.
The Pubmed dataset is a citation network of scientific publications. Each node represents a paper, and edges represent citation links. Node features are TF/IDF (Term Frequency-Inverse Document Frequency) weighted word vectors from the paper abstracts, and nodes are labeled according to their subject area.
The CiteSeer and Cora datasets are also citation networks. Nodes represent documents, and edges indicate citations between them. Node features are bag-of-word vectors, and each document is categorized into a specific topic.
Dataset Nodes Edges Classes
Texas 183 279 5
Wisconsin 251 450 5
Film 7600 26659 5
Cornell 183 277 5
Citeseer 3327 4552 6
Pubmed 19717 44324 3
Cora 2708 5278 7
Table 6: Dataset statistics, including the number of nodes, edges, and classes.
Chameleon and Squirrel, which were included in earlier studies, were excluded from our experiments because they have been removed from the official Geom-GCN repository due to maintenance or quality concerns.
All experiments were run on a computing cluster equipped with an NVIDIA H100 GPU (80GB HBM3), two Intel Xeon Platinum 8462Y+ CPUs, and 1.0 TiB of RAM. The code was implemented in PyTorch and executed using CUDA 12.4 with NVIDIA driver 550.54.15.. Our code is available at https://github.com/Layal-Bou-Hamdan/BSNN.
Hyperparameter tuning was done using Weights & Biases [5]. The full range of hyperparameters that we optimize using random search is listed in Table 7.
Hyperparameter Values
Hidden channels {8, 16, 32}
Stalk dim $d$ {2, 3, 4}
Layers {2, 3, 4, 5, 6}
Learning rate 0.02
Activation ELU
Weight decay (regular) Log-uniform $10^{-9.2},10^{-6.8}]$
Weight decay (sheaf) Log-uniform $10^{-11.0},10^{-6.8}]$
Input dropout Uniform $[0.0,0.9]$
Layer dropout Uniform $[0.0,0.9]$
Add LP / Add HP {0, 1} (separately for each)
KL term Enabled
Ensemble size 3
Orthogonalization Householder
Patience (early stopping) 200
Max training epochs 1000
Optimizer Adam
Table 7: Hyperparameter ranges for BSNN models.
To perform a hyperparameter sensitivity analysis, we used the hyperparameters listed in Table 8 with node splits of 32% for training, 20% for validation, and 48% for testing.
@@
Table 8: Hyperparameter selections for sensitivity analysis using 32%/20%/48% node splits for training, validation, and testing.